# Optimum Spectrum Extension for PAPR Reduction of DFT-s-OFDM

### v1

@@ V-C4 Rate-PAPR Trade-Off
We show that they are two distinct optima for SE size: one for PAPR minimization ($N_{\rm e}^{({\rm min\,PAPR})}$), and another for rate maximization ($N_{\rm e}^{({\rm max\,Capa})}$), which implies the existence of a range of SE sizes that can improve both PAPR and rate simultaneously. Fig. 15 compares $N_{\rm e}^{({\rm min\,PAPR})}$ and $N_{\rm e}^{({\rm max\,Capa})}$ as functions of shaping for both window types with $N_{\rm sc}=96$. For $N_{\rm e}^{({\rm max\,Capa})}$, the TDL-C (300ns) channel is considered at SNR 5-5, 0, and 5 dB. The $N_{\rm e}^{({\rm min\,PAPR})}$ curves in Figs. 15(a) and  15(b) are obtained by minimizing the bound (13) for QPSK.
@@
For both window types, the SE size optimal for PAPR minimization is larger than that for rate maximization, and the gap widens with SNR. In general, the Kaiser window requires smaller SE sizes for both criteria. Interestingly, SE improves the rate of FDSS (i.e. $N_{\rm e}^{({\rm max\,Capa})}>0$) only when the FDSS attenuation is sufficiently large.
@@
Fig. 14 also reports the achievable rate when using $N_{\rm e}^{({\rm min\,PAPR})}$ as in Figs. 15(a) and  15(b). With little shaping, this is detrimental (a loss of $\approx 5$% with no FDSS compared to plain DFT-s-OFDM), but with stronger shaping, $N_{\rm e}^{({\rm min\,PAPR})}$ can increase the rate. At 14-14dB attenuation, the rate loss with Hann window is 21% compared to plain DFT-s-OFDM (vs. 26% with no SE, and 19% with optimal SE), while for the Kaiser window it is of 17% (vs. 19% with no SE, and 13% with optimal SE).
@@
Similarly, while not optimal for PAPR, the SE size $N_{\rm e}^{({\rm max\,Capa})}$ can still provide significant PAPR reduction, as shown on Figs. 15(c) and 15(d), which give the $10^{-3}$ CCDF PAPR values using the SE size on Figs. 15(a) and  15(b). The PAPR reduction with $N_{\rm e}^{({\rm max\,Capa})}$ is nearly optimal for the Hann window at large attenuation, but weaker for the Kaiser window.
Overall, $N_{\rm e}^{({\rm max\,Capa})}$ and $N_{\rm e}^{({\rm min\,PAPR})}$ define, respectively, a lower and upper bound on the SE size to be considered for a practical PAPR–rate trade-off under a given configuration and scenario.
@@ VI Conclusion
Upper bound on $\max_{0\leq n<N_{\rm fft}}|s[n]|^{2}$ then follows by adapting [4, Th. 1] and [5, Th. 1] provided for $\phi$-BPSK and QPSK, respectively, to any constellation and shifts. The case considered therein corresponds to $L=0$, with a shift $L$ then $\angle\frac{p_{i}[n]}{p_{j}[n]}=-(i-j)2\pi\frac{L}{N_{\rm data}}+\angle\frac{p_{0}\left[n-\frac{N_{\rm fft}}{N_{\rm data}}i\right]}{p_{0}\left[n-\frac{N_{\rm fft}}{N_{\rm data}}j\right]}$ where the last term is evaluated in [4, Lem. 1].
@@
Then, as shown in [4, Lem. 1] if the FDSS window is real and symmetric we have
@@
which can also be directly verified to hold true for the Dirichlet pulse (8). Therefore we get the pulsed phase difference

### v2

@@ V-E Rate-PAPR Trade-Off
We showed that they are two distinct optima for SE size: one for PAPR minimization ($N_{\rm e}^{({\rm min\,PAPR})}$), and another for rate maximization ($N_{\rm e}^{({\rm max\,Capa})}$), which implies the existence of a range of SE sizes that can simultaneously improve both PAPR and rate. Accordingly, $N_{\rm e}^{({\rm max\,Capa})}$ and $N_{\rm e}^{({\rm min\,PAPR})}$ provides, respectively, a lower and upper bound on the SE size relevant for PAPR–rate trade-off.
@@
Fig. 19 compares $N_{\rm e}^{({\rm min\,PAPR})}$ and $N_{\rm e}^{({\rm max\,Capa})}$ as functions of the FDSS shaping for both window families with $N_{\rm sc}=96$. For $N_{\rm e}^{({\rm max\,Capa})}$, the TDL-C(300ns) channel is considered at SNR 5-5, 0, and 5 dB. The $N_{\rm e}^{({\rm min\,PAPR})}$ curves in Figs. 19(a) and  19(b) are obtained by minimizing the bound (13) for QPSK. For both window types, the optimal SE size for PAPR minimization is larger than that for rate maximization, and the gap widens with SNR. In general, the Kaiser window requires smaller SE sizes for both criteria. Interestingly, SE improves the rate of FDSS (i.e. $N_{\rm e}^{({\rm max\,Capa})}>0$) only when the FDSS attenuation is sufficiently large.
@@
Fig. 18 reports the capacity if using $N_{\rm e}^{({\rm min\,PAPR})}$ as identified in Figs. 19(a) and  19(b). With weak FDSS shaping, this choice is detrimental (e.g. a loss of $\approx 5$% with no FDSS compared to plain DFT-s-OFDM at 5 dB SNR), whereas with stronger shaping, $N_{\rm e}^{({\rm min\,PAPR})}$ can increase the rate. At 14-14dB attenuation and 5 dB SNR, the rate loss with Hann window is 21% relative to plain DFT-s-OFDM (vs. 26% with no SE, and 19% with optimal SE), while for the Kaiser window it is of 17% (vs. 19% with no SE, and 13% with optimal SE).
@@
Similarly, although not optimal for PAPR minimization, the SE size $N_{\rm e}^{({\rm max\,Capa})}$ can still provide significant PAPR reduction, as shown on Figs. 19(c) and 19(d), which report the $10^{-3}$ CCDF PAPR values using the SE size from Figs. 19(a) and  19(b). The PAPR reduction with $N_{\rm e}^{({\rm max\,Capa})}$ is nearly optimal for the Hann window at large attenuation, but weaker for the Kaiser window.
A more specific SE size for rate-PAPR trade-off could in principle be obtained by incorporating an explicit SNR loss into the capacity formula to account for the required power back-off associated with a given PAPR level. As discussed in Sec. III, the CM has been used in past 3GPP studies as a proxy for power back-off estimation, but its linear fitting used in 3GPP has not been updated since 4G, and CM exhibits a non-monotonic behavior with FDSS shaping that is unlikely to yield reliable conclusions. Alternatively, linear fittings of PAPR on power de-rating are scarcely available and would require further validation in the presence of FDSS. Moreover, a very precise optimum SE ratio is unlikely to be implementable in practice due do resources configuration constraints. In the meantime, the CM-based approach is gradually being abandoned in 3GPP in favor of more comprehensive but significantly less tractable RF power back-off simulations.
The analysis provided here helps delimit a practical range of SE sizes and provides useful rule of thumb for system design. For example, in the scenarios considered in Fig. 19 and Fig. 18, when the SNR is low, the capacity variation is small and it is preferable to select a SE size close to $N_{\rm e}^{({\rm min\,PAPR})}$, ranging from 10% to 35% depending of the FDSS window attenuation. As the SNR increases and for weak FDSS attenuation, the capacity penalty induced by SE becomes more pronounced, and smaller SE sizes on the order of 5% may be preferable. Conversely, with strong FDSS attenuation, both $N_{\rm e}^{({\rm max\,Capa})}$ and $N_{\rm e}^{({\rm min\,PAPR})}$ in the range of 20%-35% yield similar improvements in both rate and PAPR, such that further optimization in between would provide only marginal gains. Finally, it is worth noting that the analysis confirms that a $25\%$ SE, advocated as a good compromise in [31] and in 3GPP Rel.-18 study, lies well within the SE range identified in Fig. 19 for the corresponding scenario.
@@ VI Conclusion
Upper bound on $\max_{0\leq n<N_{\rm fft}}|s[n]|^{2}$ then follows by adapting [26, Th. 1] and [27, Th. 1] provided for $\phi$-BPSK and QPSK, respectively, to any constellation and shifts. The case considered therein corresponds to $L=0$, with a shift $L$ then $\angle\frac{p_{i}[n]}{p_{j}[n]}=-(i-j)2\pi\frac{L}{N_{\rm data}}+\angle\frac{p_{0}\left[n-\frac{N_{\rm fft}}{N_{\rm data}}i\right]}{p_{0}\left[n-\frac{N_{\rm fft}}{N_{\rm data}}j\right]}$ where the last term is evaluated in [26, Lem. 1].
@@
Then, as shown in [26, Lem. 1] if the FDSS window is real and symmetric we have
@@
which can also be directly verified to hold true for the Dirichlet pulse (8). Therefore we get the pulse phase difference